Question:
If a die is rolled one time, find this probability:
Of getting a number greater than 2 and an even number
Total possible outcomes = 1,2,3,4,5,6 so 6 possible outcomes
Looking to find probability of getting a number greater than 2:
There are four numbers greater than 2; (3, 4, 5, 6) so a probability of 4/6 =
2/3 chance of getting a number greater than 2
Probability of getting an even number: there are 3 even numbers out of 6
possible dice sides (2, 4, 6) so 3/6 or 1/2 chance of getting and even number.
The question asks to find probability of getting both a number greater than 2 and an even number.
So: P(> 2 and even) = P(> 2) + P(even) = 2/3 + 1/2 = 7/6

This is not the answer the solution's manual displays. What am I doing incorrectly?
Your reply is GREATLY APPRECIATED!


Answer by stanbon(75887)

The question asks to find probability of getting both a number greater than 2 and an even number.
So: P(> 2 and even) = P(> 2) + P(even) = 2/3 + 1/2 = 7/6
-------
P(> 2 AND even)
Only 4 and 6 are greater than 2 AND even.
Ans: P(>2 AND even) = 2/6 = 1/3
------------
Also:: P(>2 AND even) = P(>2)*P(even) = (4/6)(3/6) = 12/36 = 1/3
--------------
Cheers,
Stan H.
---------------